OpenStudy (anonymous):

A company makes calculators, has plants in the cities of Walton and Exton. In Walton, fixed costs are $8800 a month, and Exton's are $7000. Fabrication price for Walton stands at $6, and for Exton at $7.50. Next month, the company has to fabricate 1,500 calculators. How many should be made at each plant so that the total costs are the same?

OpenStudy (anonymous):

Cost in Walton = 8800 + 6x Cost in Exton = 7000 + 7.5y x+y=1500 Both costs must be equal: 8800+6x = 7000+7.5y Solve for x and y now.

OpenStudy (anonymous):

Thank you very much, I got to solve it on the second try.